### v1

# A Novel Theoretical Approach on Micro-Nano Robotic Networks Based on Density Matrices and Swarm Quantum Mechanics

Abstract
In a robotic swarm, parameters such as position and proximity to the target can be described in terms of probability amplitudes. This idea led to recent studies on a quantum approach to the definition of the swarm, including a block-matrix representation. Here, we propose an advancement of the idea, defining a swarm as a mixed quantum state, to be described with a density matrix, whose size does not change with the number of robots. We end the article with some directions for future research.
Keywords: quantum mechanics; swarm intelligence; multi-robot coordination
MSC (2025): 03.65.Ca, 87.85.St
1 Introduction
In the last decades, there has been a growing interest in studying, analyzing and deploying teams of robots to accomplish a wide variety of challenging tasks such as target monitoring, surveillance, environmental sensing, disaster recovery and exploration [6]. On the other hand, nature has been a source of inspiration for the development of several bio-inspired algorithms [17, 37, 26, 21]. In particular, natural phenomena such as flocks of birds, insects and oceans have since long fascinated engineers and researchers, leading to the fascinating advancements made in the development and deployment of robotic swarms [33, 15]. Specifically, researchers have made progress in understanding and imitating the behavior and communication patterns in natural swarms through swarm robots. For example, the imitation of ants led to terrestrial robotics swarms [4] and ant routing/ant colony optimization [13], the imitation of flocking bird to aerial swarms [28], the observation of fish to aquatic swarms [9], and the fascination toward the microscopic world led to microscopic [8] systems. The progressive miniaturization of the robots can lead to several applications, especially in medical technology [14, 32]. Telecommunication technologies are also quickly adapting to the micro/nano scales of devices and the related research activity is increasing progressively [12, 11], making it possible for micro/nano devices to communicate via electromagnetic fields or through molecules as information relay systems. However, the smaller the robotic system, the harder it is to control. Moreover, to control such systems, one needs to first accurately model them.
Inspired by quantum mechanics, the field of quantum computing uses quantum bits (qubit) to extend the values of the standard 0, 10,\,1 to all the real values comprised between them (the interval $[0,1])$ [36]. This makes it possible to precisely model the behavior of micro/nano systems. There are some pioneering applications of quantum computing to artificial intelligence [38, 19], robotics [7, 20], vehicle communication [35], information security [16], and swarm robotics [18, 39, 2]. However there are significant challenges in utilizing quantum algorithms in the context of swarm robotics. For example, problems concerning stability [5] and robustness [3] of the algorithms have not been well-investigated, though it has been suggested that algorithms such as quantum particle swarm optimization can lead to improvements in path-planning convergence [30]. This is due to the lack of appropriate modeling formalisms for quantum swarms. Therefore, in this article, we prepare the necessary ground for advanced control problems for quantum swarms by suggesting appropriate techniques to model them mathematically.
In quantum mechanics, density matrices were introduced for the treatment of epistemic uncertainty on systems and their noisiness. Such a description can applied, under some assumptions and suitable variable adaptation, to a swarm of robots. In [23], a description of a robotic swarm is proposed, not as a density matrix, but in terms of a nested matrix. The diagonal submatrices contain the parameters of a single robot, while the off-diagonal elements contain the information exchange between the robots. Thus, the size of the matrix depends upon the degrees of freedom of each robot, the exchanged information, and the number of robots. In particular, the size of the matrix increases with the number of robots, which constitutes a computational problem for large swarms. Though the article does not specifically concern quantum robots, the authors leverage quantum computing to quantize some information (e.g., position as well as proximity to the target) to model local interactions between the robots. Moreover,  [23] only considers only the local-to-global approach, i.e., the global swarm behavior emerged from the local pairwise interactions, but not the vice-versa.
As a solution to the problems emerging in [23], in this article we propose a different approach for modeling quantum robotic swarms. Particularly, we use the notion of density matrices, which have previously been applied for the treatment of cases of epistemic uncertainty on quantum systems, for describing a swarm of robots. Particularly, we model a quantum robotic swarm as a mixed state given by a superposition of pure states that represent individual robots, mathematically indicating them with a density matrix. By shaping the density matrix for a robotic swarm, one can develop a suitable overall dynamics for the whole swarm as an operator acting on the corresponding density matrix. It is important to note that in our approach, the size of the density matrices is rendered independent from the swarm size. In fact, it only depends on the number of degrees of freedom of the system, at the condition of treating robots as particles within a “cloud”: the swarm is thus investigated as a whole, rather than as a collection of distinguishable objects. Moreover, information of local robots can also be extracted from the density matrices through partial trace operation, thus also facilitating a global-to-local approach in the modeling of quantum swarms.
Another crucial difference of our approach with [23] is the neglection of interaction terms. In particular, while [23] considers the pair-wise interaction terms between the robots, which contribute to the size of matrices obtained for the swarm, in our paper, we can neglect these interactions. This is rooted in the fact that the density matrices already consider the effect of the interaction terms through the mixing of pure quantum states. Therefore, our approach offers a scalable yet efficient solution for modeling the information contained in the swarm without introducing any unnecessary redundancies.
Figure 1: Pictorial representation of the transition from a classic robotic swarm to a quantum-robotic swarm: in a given arena, each robot is considered in terms of a probability amplitude to be found in a certain space position.
The article is organized as follows. In Section 2 we present the necessary preliminaries from quantum computing, such as density matrices and their operations, that are useful for our results. Then, in Section 3 we propose our modeling framework and present its advantages. In Section 4 we demonstrate the application of our results through simple examples and case studies. Finally, in Section 5 we summarize our results and present some extensions and applications of our approach to be considered in the future.
2 Notations and Preliminaries
2.1 Notations
The set of real numbers are denoted by $\mathbb{R}$, and $\mathbb{R}^{n}$ denotes the $n$-dimensional real space. The set of complex numbers is denoted by $\mathbb{C}$, and a complex matrix $A\in\mathbb{C}^{m\times n}$ is of size $m\times n$. We use the standard bra-ket notation to describe a quantum state: $|\psi\rangle$ is the ket vector, a quantum pure state in a (complex) inner product vector space $V$ corresponding to an eigenstate of an operator, and $\langle\psi|$ is the bra vector, the complex conjugate of $|\psi\rangle$. Moreover, the notation $\langle\psi|\psi\rangle$ indicates the inner product, i.e., the product of bra and ket vectors, while $|\psi\rangle\langle\psi|$ denotes the outer product.
For a matrix $A$, $Tr(A)$ is the trace operation, the sum of diagonal elements of a matrix. The operator $Tr$ indicates the Trace operation, defined as the sum of the diagonal elements of the matrix, that is $\sum_{i}T_{i,i}$, which is indicated as $Tr(T):=\Sigma_{i}\langle i|T|i\rangle$ in quantum mechanics. $A_{t}$ is a generic time-dependent operator; its complex conjugate is $A_{t}^{{\dagger}}$, such that $(A^{{\dagger}}_{t})^{{\dagger}}=A_{t}$. If $A$ is Hermitian (self-adjoint), thus it is equal to its complex conjugate: $A_{t}^{{\dagger}}=A_{t}$. Finally, we use $\otimes$ to describe the tensor product.
2.2 Preliminaries
In quantum mechanics, the behavior of a quantum state is described by the Schrödinger equation [34]:
$i\hbar\frac{d}{dt}|\psi(t)\rangle=\mathbf{H}|\psi(t)\rangle,$ (1)
where $\mathbf{H}$ is the Hamiltonian operator, $i$ is the imaginary unit, $\hbar$ is the reduced Planck constant that will be set equal to 1 here, as often done in quantum physics, and $|\psi(t)\rangle$ represents the wavefunction at time t as a mathematical description of the pure quantum state in terms of its probability amplitude. When dealing with an ensemble of quantum states, such as in a quantum robotics swarm, one considers a mixture of different pure quantum states, known as mixed states. In this case, one uses the more general master equation [25, 27]
$\frac{d}{dt}\rho(t)=\mathcal{L}(\rho(t)),$ (2)
where $\rho(t)$ is the density operator, generalizing the mathematical description of the mixed quantum states, and $\mathcal{L}$ is the Lindblad operator given by
$\begin{split}&\mathcal{L}(\rho(t))=-i[H,\rho(t)]+\\ &\sum_{\alpha}\gamma_{\alpha}(t)\left(L^{T}\rho(t)L^{\ast}-\frac{1}{2}L^{{\dagger}}L\rho(t)-\frac{1}{2}\rho(t)L^{{\dagger}}L\right),\end{split}$ (3)
where the operator $L$ represents the interaction of the system with the environment. Eq. (3) corresponds to a Markovian dynamics when $\gamma_{\alpha}\geq 0$, for all $\alpha$.
[…]
Definition. A density operator [31, 27], described by a density matrix, is a Hermitian operator ($\rho(t)=\rho(t)^{{\dagger}}$), positive ($\rho(t)>0$) and with unit trace ($Tr[\rho(t)]=1$).
We now analyze some properties of the density operator and the corresponding matrix. For a pure state, the density matrix is defined as a projector
$\rho(t)=|\psi(t)\rangle\langle\psi(t)|,$ (4)
while for a mixed state we have
$\rho(t)=\Sigma_{i}p_{i}|\psi_{i}(t)\rangle\langle\psi_{i}(t)|,$ (5)
where $p_{i}$ is the $i$-th weight, $p_{i}\in\mathbb{R}$, and $\Sigma_{i}p_{i}=1$. The index $i$ denotes the $i$-th robot. It can be easily observed from the above that a pure state cannot be represented as the mixture of other states. Moreover, while a pure state is described by the wavefunction $|\psi(t)\rangle$, a mixed state is given by a superposition $\Sigma_{i}p_{i}|\psi_{i}(t)\rangle$. Note that for a pure state, we have $\rho^{2}(t)=\rho(t)$ and $Tr[\rho^{2}(t)]=1$.
As a general note: to each quantum state, one can determine a corresponding a density matrix, but given a density matrix, there can be more than a single quantum state that corresponds to it, up to a phase difference [27]. In fact, the density matrix is only a statistical object which captures the distribution of qubit information. The process of mapping a density matrix to a possible quantum state is called quantum state tomography, but it is a complex process that presents no particular advantage to us and therefore, irrelevant to this work  [27]. We present now the following remark that will be useful for our research.
[…]
Remark. Two quantum states with similar density matrices are themselves similar, even though multiple states can correspond to a single density matrix. This similarity, or closeness between the density matrices is quantified using a metric known as fidelity, with with values comprised between 0 (low fidelity) and $1$ (high fidelity) [27]. A commonly used metric for this is the trace distance, which provides information on how distinguishable two quantum states are. The trace distance between two density matrices $\rho_{1}(t)$ and $\rho_{2}(t)$ is defined as:
$D[\rho_{1}(t),\rho_{2}(t)]=\frac{1}{2}Tr|\rho_{1}(t)-\rho_{2}(t)|,$ (6)
with
$|\rho_{1}(t)-\rho_{2}(t)|\coloneq\sqrt{(\rho_{1}(t)-\rho_{2}(t))^{t}(\rho_{1}(t)-\rho_{2}(t))}.$ (7)
For a system that consists of two parts (e.g. two swarm particles, a particle and an agglomerate of particles, or two agglomerates of particles) A and B, each of them described by density matrices $\rho_{A}(t)$ and $\rho_{B}(t)$ respectively, one can describe the density matrix of the system by
$\rho_{AB}(t)=\rho_{A}(t)\otimes\rho_{B}(t).$ (8)
To recover the information on state A from the matrix $\rho_{AB}(t)$, we need to compute its partial trace with respect to the state B, which is defined as
$\begin{split}Tr_{B}[\rho_{AB}(t)]&=\Sigma_{B}[\rho_{AB}(t)]\\ &=\Sigma_{i}\rho_{A}(t)\otimes\langle i|\rho_{B}(t)|i\rangle\\ &=\rho_{A}(t),\end{split}$ (9)
where $i$ indicates the qubit which can be 0 or $1$, and thus the sum involves the elements $\langle 0|...|0\rangle$ and $\langle 1|...|1\rangle$, that is, the measurement of the density operator along states 0 and $1$, respectively. The operation of partial trace is helpful, for example, to detach the information of a system from its environment. When we have a system composed by multiple elements, we can consider the sub-system of interest as the system, and the other elements as part of the environment (as stated in the previous section).
In Section 3, we use the idea of tensor product to build up a density matrix for the robotic swarm starting from individual density matrices, and the partial trace to achieve the inverse result, i.e., recovering information on single robots from the overall matrix.
3 The Proposed Theoretical Approach: Density matrix for a robotic swarm
Let us here propose a density-matrix depiction of the swarm of robots. To do so, similarly to [23], we consider the position of each robot and its success in target finding as probability amplitudes. Specifically, these amplitudes are used the construct the density matrices of the single robots. Then, the density matrix of the swarm can be obtained via the tensor product of the individual density matrices.
Let us consider an $\mathcal{R}^{2}$ mobility plane ($x$ and $y$ axes), with a certain degree of success in reaching the target. Let $|\psi_{R_{i}}(t)\rangle$ be the state of robot $R_{i}$ at time $t$, defined as
$\begin{split}|\psi_{R_{i}}(t)\rangle=&\alpha_{i}^{1}(t)|111\rangle+\alpha_{i}^{2}(t)|110\rangle+\alpha_{i}^{3}(t)|100\rangle+\\ &\alpha_{i}^{4}(t)|000\rangle+\alpha_{i}^{5}(t)|001\rangle+\alpha_{i}^{6}(t)|010\rangle+\\ &\alpha_{i}^{7}(t)|101\rangle+\alpha_{i}^{8}(t)|011\rangle.\end{split}$ (10)
$|\psi_{R_{i}}(t)\rangle$ is the superimposition state vector of 3 qubits, of the form $|a,b,c\rangle$, with qubits a,b,ca,\,b,\,c, where the first qubit indicates the position projection along a segment $[0,1]$ on the x-axis (0 and $1$ are the two extremes), the second qubit denotes the projection of the position along a segment $[0,1]$ on the y-axis (similarly defined); the coefficients $\alpha_{i}^{j}(t)$, with $i$ indicating the robot, and $j$ indicating the eigenvector, denote the probability amplitudes for the respective eigenstates/eigenvectors, satisfying $\sum_{j=1}^{8}|\alpha_{i}^{j}(t)|^{2}=1$. Then, for a robot $R_{i}$, the density operator is thus
$\rho_{R_{i}}(t)=|\psi_{Ri}(t)\rangle\langle\psi_{R_{i}}(t)|,$ (11)
and the state of $R_{i}$ can be written as a column matrix
$|\psi_{R_{i}}(t)\rangle=\left(\begin{matrix}\alpha_{i}^{1}(t)\\ \alpha_{i}^{2}(t)\\ \dots\\ \alpha_{i}^{8}(t)\\ \end{matrix}\right).$ (12)
The density matrix is thus obtained via the product between the column matrix and the row matrix of complex conjugates coefficients, as follows:
$\rho_{R_{i}}(t)=\left(\begin{matrix}\alpha_{i}^{1}(t)\\ \alpha_{i}^{2}(t)\\ \vdots\\ \alpha_{i}^{8}(t)\\ \end{matrix}\right)\left(\alpha_{i}^{1,\ast}(t),\;\alpha_{i}^{2,\ast}(t),\;\dots,\;\alpha_{i}^{8,\ast}(t)\right).$ (13)
[…]
Remark. In this article, we only consider real values for the probability amplitudes, so the complex conjugates are equal to the original values. The time evolution of a single robot can be obtained via a suitable time-evolution operator acting on the corresponding density operator. Such an operator can include a dependence on the state of the other robots, acting according to the scope of the robotic mission. In [23], the swarm behavior emerged from pairwise robotic interactions modeled upon a quantum circuit. Here, we consider a more quantum-mechanics style approach. Also, the matrix size does not increase when the swarm becomes bigger (with more units).
Then, the density matrix of a swarm of $N$ robots is computed as
$\rho_{swarm}(t)=\rho_{R_{1}}(t)\otimes...\otimes\rho_{R_{N}}(t),$ (14)
and, the information on the $i^{\text{th}}$ robot can be retrieved performing the partial trace on all the other robots using
$\rho_{R_{i}}(t)=Tr_{R_{j\neq i}(t)}(\rho_{swarm}(t)),$ (15)
where $Tr_{R_{j\neq i}(t)}$ denotes the sum of the diagonal elements of $\rho(t)$ with the exclusion of the ones indexed as $i$, that is, corresponding to the $i$-th robot. The formulation of a quantum swarm in terms of tensor products of density matrices unfortunately has a major drawback: it results in large matrix sizes for large swarm sizes. An alternate approach to alleviate this problem is to consider the swarm as a mixed state given by the superposition of pure states that represent single robots. This reduces the size of the matrix associated with the robot swarm, making the description more concise, compared to [23]. Indeed, one can also consider single robots as mixed states, in which case, one needs to resort once again to consider the tensor product of density matrices for a clear representation of density matrix of the swarm.
To illustrate this, consider an $N$-robot swarm, where the density matrix of the $i^{\text{th}}$ robot is given by $\rho_{i}(t)=|\psi_{R_{i}}(t)\rangle\langle\psi_{R_{i}}(t)|$. Considering the swarm as a mixed state, its density matrix can be simply computed as
$\rho_{swarm}(t)=\sum_{i=1}^{N}w_{i}\rho_{i}(t),$ (16)
where $w_{i}$, with $i\in\{1,\ldots,N\}$ robots, are constant weights such that $\sum_{i}w_{i}=1$. If the contribution of all the robots is equal, eq. (16) becomes
$\begin{split}\rho_{swarm}(t)&=\frac{1}{N}\sum_{i=1}^{N}\rho_{i}(t)=\\ &=\frac{1}{N}\sum_{i=1}^{N}|\psi_{R_{i}}(t)\rangle\langle\psi_{R_{i}}(t)|\end{split}$ (17)
Eq. (16) involves the sum of matrices, not their tensor product, and thus the dimension of the density matrix of the swarm, while seen as a mixed state, remains the same, as the dimensions of the single density matrices for the robots.
Moreover, while the swarm modeling approach presented in [23] utilizes interaction terms in the swarm submatrices to describe the interactions between robots, in this work, we neglect the interaction terms by considering that each robot sends all information to all the other robots, and each robot receives information from all the other robots. Thus, copying this information in the interaction sub-matrices would just be redundant. Note that the effect of “interaction” are automatically included in the density-matrix depiction of the swarm, seen as a whole system in a mixed state, resulting from the mixing of single states of the single robots. However, the utilization of explicit interation terms largely depends on the size of the considered devices: while bigger robots may use classical wireless communications, for micro/nano scale devices, newer opportunities arise [1], since they are nearer to quantum systems [29].
In the following section, we illustrate our proposed method with several examples.
4 Examples
4.1 Pure States
As a first example, we consider a toy-swarm of two robots, $R_{1}$ and $R_{2}$, moving along a line of length $1$, i.e., from $x=0$ to $x=1$. Suppose that the target is located in $x=1$. From Eq. (10), we can write
$|\psi_{R_{i}}(t)\rangle=\alpha_{i}^{1}(t)|11\rangle+\alpha_{i}^{2}(t)|10\rangle+\alpha_{i}^{3}(t)|01\rangle+\alpha_{i}^{4}(t)|00\rangle,$ (18)
where the first qubit indicates the position along the $x$ axis, and the second qubit the degree of success in target finding, and $\sum_{j}^{4}|\alpha_{i}^{j}(t)|^{2}=1$. With a slight abuse of notation, we represent the position of the robot instead with the peak amplitude of the wavefunction of the position [23]. Moreover, we remove the time argument from the state for notational simplicity by assuming that the positions are considered at a pre-defined time instant. Now suppose that $R_{1}$ is at along $x=0.5$ (the peak of the wavefunction is centered in $0.5$), and it has 0 success (the peak of the “success” wavefunction is in 0). On the other hand, $R_{2}$ at $x=1$ and with a successful result in target finding ($1$); it is thus in an eigenstate. Substituting the non-null coefficients $\alpha^{2}_{1}=\alpha^{4}_{1}=\sqrt{1/2}$ and $\alpha^{1}_{2}=1$ for the two robots, respectively, we obtain
$|\psi_{1}\rangle$ $=\frac{1}{\sqrt{2}}(|10\rangle+|00\rangle),\,$ (19)
$|\psi_{2}\rangle$ $=|11\rangle,$ (20)
Correspondingly, the density matrix of $R_{1}$ is given by
$\rho_{1}=|\psi_{1}\rangle\langle\psi_{1}|=\frac{1}{2}(|10\rangle\langle 10|+|10\rangle\langle 00|+|00\rangle\langle 10|+|00\rangle\langle 00|),$ (21)
and that of $R_{2}$ is given by
$\rho_{2}=|\psi_{2}\rangle\langle\psi_{2}|=|11\rangle\langle 11|.$ (22)
In matrix form, we have
$\rho_{1}=\frac{1}{2}\left(\begin{matrix}0&0&0&0\\ 0&1&0&1\\ 0&0&0&0\\ 0&1&0&1\\ \end{matrix}\right),\,\rho_{2}=\left(\begin{matrix}1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ \end{matrix}\right).$
The density matrix of the toy swarm is obtained as their tensor product. If we have the state as the initial information, and we want to recover the probability of the robot to be in $x=1$, we compute:
$|\langle 1_{x}|\psi_{1}\rangle|^{2}=0.5,\,|\langle 1_{x}|\psi_{2}\rangle|^{2}=1.$ (23)
The probability amplitude that the swarm barycenter is in $x=1$ is the mean of the two terms in Eq. (23), i.e., $0.75$. An ideal swarm with two robots, both centered on the target in $1$ and successful, of course has a barycenter with probability $1$ to be found on $1$. A measure of distance between the swarm barycenter and the ideal-swarm barycenter is the difference |0.751||0.75-1|. The information on positions of single robots can also be retrieved from the density matrix; this is useful in cases where we only have access to the density matrix. The partial trace of the individual density matrices with respect to the second qubit (taking away the information on “success”, in short $s$) provides information on the individual robot’s position. Concerning $R_{1}$, after a few passages we obtain:
$\begin{split}&Tr_{s}\rho_{1}=\langle 1_{s}|\rho_{1}|1_{s}\rangle+\langle 0_{s}|\rho_{1}|0_{s}\rangle=\\ &=\frac{1}{2}(|1\rangle\langle 0|+|0\rangle\langle 0|),\end{split}$ (24)
while for $R_{2}$ we simply have
$Tr_{s}\rho_{2}=\langle 1_{s}|\rho_{2}|1_{s}\rangle+\langle 0_{s}|\rho_{2}|0_{s}\rangle=|1\rangle\langle 1|.$ (25)
We notice that $R_{1}$ has $1/2$ probability of being in $1$, while $R_{2}$ has probability $1$. Thus, the barycenter of the swarm is their mean, $0.75$. In this way, we prove that the same information can be recovered from density matrices. Tracing the overall density matrix $\rho_{swarm}=\rho_{12}=\rho_{1}\otimes\rho_{2}$ with respect to the second qubit, we can also recover the position of the swarm barycenter. The same applies to the ideal swarm centered on the target. The comparison between the values for the given swarm and the ideal one, gives a measure of the precision with which the target has been reached. We can define an opportune distance between matrices as the trace of the difference between the reduced matrices, that is $Tr(|\langle 1_{s}|\rho_{12}|1_{s}\rangle|-|\langle 1_{s}|\rho_{12(T)}|1_{s}\rangle|)$, where we focus on the probability of finding the barycenter close to $x=1$, and where $\rho_{12}$ is the density matrix of the ideal swarm. It is clear that we can think of a whole class of “successful” swarms, whose robots are inside a small neighbor of the target. In addition, using our modeling framework, one may also consider stability problems. Specifically, one may define a Lyapunov as the Hamiltonian of the barycenter of the swarm, which is a function of the robot’s position displacements (potential term) and momentum (speed term). Stability will be investigated in future research.
4.2 Mixed States
Consider again the 2-robot swarm described above. As discussed previously, one can significantly benefit in formalizing the robot swarm as a mixed state consisting of robots that are described as pure states. In this case, the density matrix of the swarm would be defined as the weighted sum of the states, rather than as their tensor product.
$\rho_{swarm}=\frac{1}{2}|\psi_{1}\rangle\langle\psi_{1}|+\frac{1}{2}|\psi_{2}\rangle\langle\psi_{2}|$ (26)
The information on a single robot can still be recovered via the partial trace, tracing the $\rho_{swarm}$ with respect to the other robot(s). In matrix form, Eq. (26) would give
$\rho_{swarm}=\left(\begin{matrix}1/2&0&0&0\\ 0&1/4&0&1/4\\ 0&0&0&0\\ 0&1/4&0&1/4\\ \end{matrix}\right).$ (27)
We verify that $Tr(\rho_{swarm})=1$; to check that $\rho_{swarm}$ is a mixed state, we compute the trace of the squared matrix, verifying that it is smaller than 1
$Tr(\rho_{swarm}^{2})=Tr\left(\begin{matrix}1/4&0&0&0\\ 0&1/8&0&1/8\\ 0&0&0&0\\ 0&1/8&0&1/8\\ \end{matrix}\right)=\frac{1}{2}.$
For the swarm we compute the overall energy $E$ as $E=\Sigma_{i=1}^{N}E_{i}$, with $N=2$ in our example, and $E_{i}$ the energy of each single robot. Then, we compute $\Delta E=E(t)-E(t^{\prime})$, where the time dependence is related with the instantaneous peak position, and we can finally choose the Hamiltonian as
$H=\frac{\Delta E}{2}\sigma_{z}=\frac{\Delta E}{2}(|0\rangle\langle 0|-|1\rangle\langle 1|).$ (28)
To reduce the dimensions of the problem, we can perform the partial trace operation on $\rho_{swarm}$ with respect to the success information. We also perform this operation beforehand on the density matrices of single robots, as showed in Eq. (24, 25). The reduced density matrix of the swarm as a mixed state is thus
$\begin{split}\rho_{swarm}^{r}&=\frac{1}{2}Tr_{s}(\rho_{1})+\frac{1}{2}Tr_{s}(\rho_{2})=\\ &=\frac{1}{4}(|1\rangle\langle 0|+|0\rangle\langle 0|+2|1\rangle\langle 1|).\end{split}$ (29)
If we consider the ideal situation where the interaction of the swarm with the environment is absent or negligible, the $L$ operators of the master equation in Lindblad form are negligible as well, and the master equation is reduced to
$\mathcal{L}(\rho_{swarm}^{r})=-i[H,\rho_{swarm}^{r}]=-i\frac{\Delta E}{2}(|0\rangle\langle 0|-|1\rangle\langle 1|).$ (30)
To consider the interaction with the environment, we have to include non-null $L$ operators. For the sake of simplicity, we can choose $L=|0\rangle\langle 0|$. Thus, the remaining terms of the master equation are computed as
$\begin{split}L^{T}\rho_{swarm}^{r}L^{\ast}=&\frac{1}{4}|0\rangle\langle 0|\\ -\frac{1}{2}L^{{\dagger}}L\rho_{swarm}^{r}=&-\frac{1}{8}|0\rangle\langle 0|\\ -\frac{1}{2}\rho_{swarm}^{r}L^{{\dagger}}L=&-\frac{1}{8}(|1\rangle\langle 0|+|0\rangle\langle 0|).\end{split}$ (31)
Summing up the terms of Eqq. (30,31), we obtain
$\mathcal{L}(\rho_{swarm}^{r})=-i\frac{\Delta E}{2}(|0\rangle\langle 0|-|1\rangle\langle 1|)-\frac{1}{8}|1\rangle\langle 0|.$ (32)
4.3 Other computational examples
In the next few examples, we consider another toy swarm, constituted by two robotic fish $R_{1},\,R_{2}$, moving along the segment $[0,1]$ of the $x$ axis, as shown in Figure 2. For the sake of simplicity, we neglect here information about success in target reaching, focusing only on information about position. Case A presents two robots maintaining a constant distance between them, while case B presents two robots getting distant from each other.
Let us focus on case A, an example of swarm compactness over time. Figure 2 exemplifies the passage to a quantum representation, where the positions are substituted by the peaks of the wavefunctions. As an example, we choose to write the density operator of the swarm as a mixed state:
$\rho^{A}(t_{0})=\frac{1}{2}|0\rangle\langle 0|+\frac{1}{2}(\sqrt{0.8}|0\rangle+\sqrt{0.2}|1\rangle)(\sqrt{0.8}\langle 0|+\sqrt{0.2}\langle 1|).$ (33)
The swarm at the following time point is represented by the density operator:
$\rho^{A}(t_{1})=\frac{1}{2}|1\rangle\langle 1|+\frac{1}{2}(\sqrt{0.8}|1\rangle+\sqrt{0.2}|0\rangle)(\sqrt{0.8}\langle 1|+\sqrt{0.2}\langle 0|).$ (34)
The respective matrices, for the swarm at $t_{0}$ and at $t_{1}$, are:
$\rho^{A}(t_{0})=\left(\begin{matrix}0.9&0.2\\ 0.2&0.1\end{matrix}\right),\,\,\,\,\rho^{A}(t_{1})=\left(\begin{matrix}0.1&0.2\\ 0.2&0.9\end{matrix}\right).$ (35)
To obtain these results, we have to multiply bra-kets, and sum the terms:
$\begin{split}\rho^{A}(t_{0})=\frac{1}{2}|0\rangle\langle 0|+\frac{1}{2}(\sqrt{0.8}|0\rangle+\sqrt{0.2}|1\rangle)(\sqrt{0.8}\langle 0|+\sqrt{0.2}\langle 1|)=\\ \frac{1}{2}|0\rangle\langle 0|+\left(\frac{0.8}{2}|0\rangle\langle 0|+\frac{\sqrt{1.6}}{2}|0\rangle\langle 1|+\frac{\sqrt{1.6}}{2}|1\rangle\langle 0|+\frac{0.2}{0}|1\rangle\langle 1|\right)=\\ 0.9|0\rangle\langle 0|+0.2|0\rangle\langle 1|+0.2|1\rangle\langle 0|+0.1|1\rangle\langle 1|,\end{split}$ (36)
and then, organizing the coefficients 0.9, 0.2, 0.2, 0.10.9,\,0.2,\,0.2,\,0.1 inside the matrix with columns $|0\rangle,\,|1\rangle$ and rows $\langle 0|,\,\langle 1|$, we get the result.
Supposing that the variation of energy is positive, we obtain:
$Tr[\mathcal{L}(\rho^{A}_{t_{1}})\Delta\rho^{A}]=-0.16i\Delta E<0.$ (37)
Recalling the definition of $\rho$-based stability in [10], in the case of positive $\Delta E$, the condition of Eq. (37) denotes stability for the case $A$. In this article, we chose to just present a general approach to a quantum swarm, but this formalism can be helpful for future technological developments of the research.
We conclude the section of Examples with a case of target-reaching application. Let us start from the previously-found swarm density matrix
$\rho^{S}=\left(\begin{matrix}0.9&0.2\\ 0.2&0.1\end{matrix}\right),$ (38)
and let us consider a target in $|0\rangle\langle 0|$, as an ideal swarm centered on the target, described by a state $\rho^{T}$ having density matrix
$\rho^{T}=\left(\begin{matrix}1&0\\ 0&0\end{matrix}\right).$ (39)
The distance between the swarm and the target can be computed with the trace distance definition:
$\begin{split}d(\rho_{S},\rho_{T})=\frac{1}{2}Tr[\sqrt{(\rho_{S},\rho_{T})^{t}(\rho_{S},\rho_{T})}]=0.05.\end{split}$ (40)
The adoption of density matrices for swarms of quantum robots can pave the way toward applications in control theory, adapting to our cases criteria of stability for open quantum systems, that are based on density matrices and suitable choices of Lyapunov functions [10].
Figure 2: Quantum-like representation of a toy 2-robot swarm, where the position of the robot is considered as the weighted peak of the corresponding wavefunction.
4.4 Time evolution and enhanced representations
We propose here a working example to visualize a toy swarm of two robots as surfaces representing the wavefunctions. We also sketch a procedure to compute the time evolution from one swarm configuration to another one. The process describes a global-local approach, wherein density matrices of the swarm and a time evolution operator are computed to determine the evolution of the swarm. Then, local distribution of the individual robots are updated accordingly. Such a process, as opposed to the local-to-global approach proposed in [23], allows to streamline local information of the robots from global information on the swarm. Suppose that the initial state of particle 1 at time $t_{0}$ is
$|\psi_{1}(t_{0})\rangle=\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle),$ (41)
and the initial state of particle 2 is
$|\psi_{2}(t_{0})\rangle=|00\rangle$ (42)
Thus, the density matrix for the first particle-robot is
$\rho_{1}(t_{0})=\frac{1}{2}(|00\rangle\langle 00|+|11\rangle\langle 11|)=\frac{1}{2}\left(\begin{matrix}1&0&0&1\\ 0&0&0&0\\ 0&0&0&0\\ 1&0&0&1\end{matrix}\right),$ (43)
and the density matrix for the second particle-robot is
$\rho_{2}(t_{0})=|00\rangle\langle 00|=\left(\begin{matrix}1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{matrix}\right).$ (44)
From this, we obtain the mixed density matrix as
$\rho_{mix}(t_{0})=\frac{\rho_{1}(t_{0})+\rho_{2}(t_{0})}{2}=\left(\begin{matrix}0.75&0.&0.&0.25\\ 0.&0.&0.&0.\\ 0.&0.&0.&0.\\ 0.25&0.&0.&0.25\end{matrix}\right)$ (45)
We choose the final state of particle 1 at $t_{1}$ as
$|\psi_{1}(t_{1})\rangle=\frac{\sqrt{0.2}}{\sqrt{0.8^{2}+0.2^{2}}}|00\rangle+\frac{\sqrt{0.8}}{\sqrt{0.8^{2}+0.2^{2}}}|11\rangle,$ (46)
and that of particle 2 as
$|\psi_{2}(t_{1})\rangle=\frac{\sqrt{0.1}}{\sqrt{0.9^{2}+0.1^{2}}}|00\rangle+\frac{\sqrt{0.9}}{\sqrt{0.9^{2}+0.1^{2}}}|11\rangle.$ (47)
Finally, the density matrix of the mixed state is given by
$\rho_{mix}(t_{1})=\left(\begin{matrix}0.15&0.&0.&0.35\\ 0.&0.&0.&0.\\ 0.&0.&0.&0.\\ 0.35&0.&0.&0.85\end{matrix}\right).$ (48)
Figure 3 shows the robots corresponding to the two “particles” as surfaces, whose height along the $z$-axis denotes the probability amplitude. Such a representation is an additional overcoming of the visualizations of [23], where a classic position was considered in correspondence to the peak of the wavefunctions. These visualizations have been obtained through Python - Jupyter Notebook.
Figure 3: Surface visualization of the probability amplitudes to find the two robots-particles described in Eq. (45) and (48), respectively.
We are also interested in finding a method to approximate the time evolution operator, that is, a unitary matrix $O(t)$ such that:
$O(t_{1})\rho_{mix}(t_{0})=\rho_{mix}(t_{1}).$ (49)
A possible technique is the singular value decomposition (SVD). We first compute the SVD of $\rho_{mix}(t_{0})$ obtaining two matrices $U$ and $V^{{\dagger}}$ and then the SVD of $\rho_{mix}(t_{1})$ obtaining two other matrices $U^{\prime}$ and $V^{\prime{\dagger}}$, such as:
$\rho_{mix}(t_{0})=USV^{{\dagger}},\,\,\rho_{mix}(t_{1})=U^{\prime}S^{\prime}V^{\prime{\dagger}},$ (50)
where $S$ (respectively $S^{\prime}$) is a diagonal matrix with the singular values of $\rho_{mix}(t_{0})$ (respectively $\rho_{mix}(t_{1})$). These new matrices can be used to reconstruct the process from $\rho_{mix}(t_{0})$ to $\rho_{mix}(t_{1})$. In fact, the needed time-evolution operator is:
$O=U^{\prime}(V^{\prime{\dagger}})^{-1}VU^{{\dagger}}.$ (51)
The choice of the dynamics is fundamental to shape the global motion of a swarm. This can be useful in a search-and-rescue scenario, when the robots exchange information about individual target proximity, and then the overall swarm reaches a point in space. The point in space is represented as an ideal swarm centered on it. In our case, it can be the target density matrix. The approximation of the time operator represented the required overall transformations to let the swarm approximately reach that region. We can also apply to $\rho_{mix}(t_{0})$ some given specific dynamics, and observe how the swarm is moving accordingly. A classic example is $U=e^{-iHt}$. The systematic definition of density matrices for a swarm, and a set of subsequent computations of the dynamics between pairs of time points, can allow one to approximate the functions inside the dynamics. Such an approach can foster precise applications in the domain of global-to-local control of the swarms.
In particular, we can define the algorithm illustrated in Figure 4. Our formalism can thus give some theoretical prediction on target reaching that can be experimentally verified. Let us describe the idea of Figure 4. We start with the local dimension, defining an initial distribution of robots in terms of $|\psi_{i}\rangle$. Then we compute their density matrix distribution, thus going into the Global dimension. We also need a target position defined as an estimated probability density (expressed inside a $\rho_{T}$), which is estimated, and thus we call it $\rho_{T}^{e}$. This information will be progressively refined. At this point we can approximate a dynamics $U$ such as $U\rho_{swarm}\sim\rho_{T}^{e}$. We can update the swarm distribution in terms of density matrix. And we also have to update the distribution of robots at the local dimension. Since more individual configuration correspond to the same density matrix distribution, the robots can be shifted with the smallest displacements such that their $\psi_{i}$ approximate the new $\rho_{swarm}$. The robots activate their sensors, and detect their individual proximity to the target. This information allow us to update the target distribution $\rho_{T}^{e}$, which allows for a new dynamics computation. The cycle continues until the swarm reaches the target with a great precision. This method, inspired by optimization techniques, can help yield prediction on the number of passages for target reaching.
Figure 4: An algorithm for swarm-robotic target reaching based on the defined matrices, and including an interaction between local and global.
5 Conclusions and Discussions
In this article, we proposed a new modeling methodology for a quantum robotic swarm using density matrices to describe swarms of micro and nano robots under quantum effects in a compact yet precise manner. We also proposed some toy examples illustrating applications of our modeling formalism. The computations are straightforward and have been performed by hand. More complex computations including the computation of density matrices may also be performed via the open-source software Qiskit, [^1] https://www.ibm.com/quantum/qiskit automatizing the approach.
In the future, one may consider the application of our approach to describe the motion of single robots as well as the center of mass of the quantum swarm. Experiments on physical robots may also help investigate the limits of our approach, especially concerning the necessary calibration of parameters representing divergence and convergence, which also depend upon the size of the swarm, the amplitude of movements of their components. Future research can focus on the matrix definition of multiple swarms and hierarchical approaches. In fact, we can define the density matrix of a system with multiple swarms as the tensor product of the density matrices of each swarm, or as the density matrix of the whole swarm seen as a mixed state of individual swarms. The information on the single swarm can vice versa be recovered performing the operation of partial trace with respect to all the other swarms, considered as “environment,” to the overall density matrix.
Further studies can exploit the density-matrix based formalism to develop control approaches for quantum robots [10]. In fact, our formalism with density matrices for robotics swarms can pave the way toward future applications of stability criteria to the robotics swarms, adapting control criteria defined for open quantum systems to the robots [10]. Moreover, one can also address the matrix depiction of multi-layer systems, recently used to characterize hierarchical systems of swarms communicating with mobile computing centrals [22], whose information-sharing systems are shaped according to the protocols of telecommunications [24].
Funding
The work by M. A. is funded by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/1 – Project ID 390696704 – Cluster of Excellence “Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of Technische Universität Dresden.